# Quasinormal modes and bound states of massive scalar fields in generalized Bronnikov-Ellis wormholes

### 05 Mar 2024 v2

@@ Abstract
In this work we explore generalized Bronnikov-Ellis wormholes backgrounds and investigate the propagation of massive scalar fields. The first generalized wormhole geometry arises from General Relativity with a phantom scalar field while the second one arises from $f(R)$-gravity. For the first geometry we employ the continued fraction method to calculate accurately the quasinormal frequencies of massive scalar fields, particularly focusing on low values of the angular number, and we show an anomalous behaviour of the decay rate of the quasinormal frequencies, for $n\geq\ell$. Also, we show that for massive scalar field the effective potential allows potential wells for some values of the parameters which support bound states which are obtained using the continued fraction method and these are characterized by having only a frequency of oscillation and they do not decay. For the second wormhole geometry the quasinormal modes can be obtained analytically, being the longest-lived modes the one with lowest angular number $\ell$. So, in this wormhole background the anomalous behaviour is avoided.
@@ I Introduction
Solutions of Einstein equations that connect different parts of the Universe or two different Universes can be found in General Relativity (GR) known as wormholes. The concept of the wormhole can be traced back to Flamm in 1916 Flamm and then wormhole-type solutions were found by Einstein and Rosen (ER), where they considered a physical space representing a physical space which was connected by a wormhole-type solution known as ER bridge Rosen . Then the pioneering work of Misner and Wheeler MisWheel and Wheeler Wheel developed further the wormholes concept. Lorentzian wormholes were studied in General Relativity (GR) Bronnikov:1973fh ; Ellis:1973yv ; Morris ; phantom ; Visser , in which conditions for traversable wormholes were found introducing a static spherically symmetric metric. However, a condition on the wormhole throat leads to the violation of the null energy condition (NEC). To allow the formation of traversable wormhole geometries a matter distribution of exotic or phantom matter has to be introduced phantom .
Recently there are many efforts to understand modifications of GR which are generated by the presence of a scalar field coupled to gravity. This coupling has important implications in local and global solutions in these theories known as scalar-tensor theories Fujii:2003pa . One of the best studied scalar-tensor theories is provided by the Horndeski Lagrangian Horndeski:1974wa . The Horndeski theory has been studied in short and large distances. In this theory black hole solutions were found Kolyvaris:2011fk ; Rinaldi:2012vy ; Kolyvaris:2013zfa ; Babichev:2013cya ; Charmousis:2014zaa , known as Galilean black holes, and also wormhole geometries were generated Korolev:2014hwa -Bakopoulos:2023tso
@@
An important issue of the compact objects formed by black holes or wormholes is their stability. The relativistic collision of two of these compact objects produces gravitational waves which will help us to understand better the gravitational interaction. However, the recent LIGO detections Abbott:2016blz -TheLIGOScientific:2017qsa do not yet probe the detailed structure of spacetime beyond the photon sphere. Future observations may detect the ringdown phase, which is governed by a series of damped oscillatory modes at early times, which are known as quasinormal modes (QNMs) Vishveshwara:1970zz -Cardoso:2016rao Future Gravitational observations may provide information on different compact objects than black holes, objects without event horizons which are known as exotic compact objects (ECOs) Mazur:2001fv -Holdom:2016nek . Perturbations of black holes and wormholes were extensively studied in the literature. In the case that a test scalar perturbs black hole compact objects the partially reflected waves from the photon sphere (PS) mirror off the AdS boundary and re-perturb the PS to give rise to a damped beating pattern. In the case of a wormhole do not decay with time but have constant and equal amplitude to that of the initial ringdown. Therefore the knowledge of QNMs and quasinormal frequencies (QNFs) can give us important information on the properties of compact objects and distinguish their nature. The QNMs give an infinite discrete spectrum which consists of complex frequencies, $\omega=\omega_{R}+i\omega_{I}$, where the real part $\omega_{R}$ determines the oscillation timescale of the modes, while the complex part $\omega_{I}$ determines their exponential decaying timescale (for a review on QNM modes see Kokkotas:1999bd ; Berti:2009kk ).
@@
QNMs and QNFs have also been calculated in the background of wormholes. The perturbations of wormhole spacetime was investigated in Konoplya:2005et . It was found that the QNMs behave the same way as in black hole background and includes the quasinormal ringing and power-law asymptotically late time tails. However, in Konoplya:2016hmd it was found that the symmetric WHs do not have the same ringing behaviour of the BHs at a few various dominant multipoles. In the case of massive probe scalar field perturbations, it was shown Churilova:2019qph that wormholes with a constant redshift function do not allow for longest-lived modes, while wormholes with non-vanishing radial tidal force do allow for quasi-resonances. Calculating the high frequency QNMs modes in the case of the Morris-Thorne wormhole spacetime, the shape function of a spherically symmetric traversable Lorenzian wormhole near its throat was constructed in Konoplya:2018ala . Scalar and axial QNMs of massive static phantom WHs were calculated in Blazquez-Salcedo:2018ipc .
@@
As we already discussed in the case of the Horndeski scalar-tensor theory if matter, parameterized by a phantom scalar field, is strongly coupled to gravity then bound states are formed Koutsoumbas:2018gbd ; Chatzifotis:2020oqr . In this work we follow a much simpler approach. We consider the Bronnikov-Ellis wormhole Bronnikov:1973fh ; Ellis:1973yv and we introduce a gravitational mass. We perturb the massive wormhole background by a test massive scalar field. Using the continued fraction method we calculate accurately the QNFs for low values of the angular number. We find stability of the background wormhole and calculating the effective potential we show that for an interplay of the mass parameter and the phantom scalar charge, using the continued fraction method potential wells are generated supporting bound states.
We also studied the QNFs in a wormhole spacetime supported by a phantom scalar field in which the scalar curvature is generalized to an $f(R)$ function Karakasis:2021lnq ; Karakasis:2021rpn ; Tang:2019qiy ; DeFalco:2021ksd ; Karakasis:2021tqx . In this model the scalar field self-interacts with a mass term potential which is derived from the scalar equation and in the resulting $f(R)$ model the scalar curvature is modified by the presence of the scalar field and it is free of ghosts and avoids the tachyonic instability.
The work is organized as follows. In Section II we review the solutions of the traversable wormholes. In Section III we study the perturbations of massive scalar field in the wormhole geometry. In Section IV we calculate the QNFs of a generalized Bronnikov-Ellis wormhole. In Section V we calculate the QNFs of a $f(R)$ gravity wormhole. In Section VI are our conclusions.
@@ II Review of traversable wormholes
The ensure the absence of horizons and singularities requires $\Phi(r)\neq 0$ which means that $\Phi(r)$ is finite throughout the spacetime. The Ellis Ellis:1973yv is a wormhole solution of an action that consists of a pure Einstein-Hilbert term and a scalar field with negative kinetic energy
@@
Given the metric ansatz (5) and setting $f(r)=R(r)$ and $V(r)=0$, the aforementioned equations yield the following solution
@@
this metric corresponds to the Bronnikov-Ellis wormhole in the limit $M\rightarrow 0$ Bronnikov:1973fh ; Ellis:1973yv . Note that the parameter $b_{0}$ specifies the charge of the phantom scalar field.
@@
For the generalized Bronnikov-Ellis wormhole, the obtained expressions are as follows
@@
III Massive scalar fields perturbations
@@
$\Psi(t,r,\theta,\phi)=e^{-i\omega t}\>\frac{R(r)}{r}\>Y_{\ell}(\Omega)\,,$ (15)
@@
$\frac{\mathrm{d}^{2}R}{\mathrm{d}r^{\ast}{}^{2}}+[\omega^{2}-V(r^{\ast})]R=0\,,$ (16)
@@
IV Generalized Bronnikov-Ellis wormhole
@@
which we plot in Fig. 1, where we can observe that have the shape of a barrier, and also can develop a well near the throat for a massive scalar field and low values of $\ell$, while for high values of $\ell$ the well disappears, so, it can support bound states for low values of $\ell$.
Figure 1: Effective potential of the generalized Bronnikov-Ellis wormhole, for $M/b_{0}=0.25$, $mb_{0}=0.8$ and $\ell=0,1,2$.
@@
$(2M-r)\left(b_{0}^{2}-r^{2}\right)R^{\prime\prime}(r)+\frac{\left(-3Mb_{0}^{2}+Mr^{2}+b_{0}^{2}r\right)}{r}R^{\prime}(r)+$ (21)
$\frac{M\left(\left(2m^{2}r^{2}+2\ell^{2}+2\ell-1\right)r^{2}+3b_{0}^{2}\right)-r\left(r^{2}\left(\ell^{2}+\ell+r^{2}(m^{2}-\omega^{2})\right)+b_{0}^{2}\right)}{r^{2}}R(r)=0\,,$
which we will solve numerically in order to obtain the QNFs.
@@ IV.1 Quasinormal frequencies
In this section we use the continued fraction method (CFM) to calculate accurately the QNFs for massive scalar fields propagating in the generalized Bronnikov-Ellis wormhole. The CFM was devised by Leaver to compute the QNMs of Schwarzschild and Kerr black holes Leaver:1985ax ; Leaver:1990zz , and improved later by Nollert Nollert:1993zz . It has been used to computed the QNMs in several black holes backgrounds, see for instance Cardoso:2003qd ; Konoplya:2013rxa ; Richartz:2014jla ; Chowdhury:2018pre ; Becar:2022wcj . However, to our knowledge this is the first time that this method is applied to calculate the QNFs in wormhole geometries.
To solve the radial equation it is necessary to specify the boundary conditions. Given the asymptotically flat nature of the wormhole geometry and the fact that the effective potential forms a barrier approaching $V(r^{\ast})\rightarrow m^{2}$ as $r^{\ast}\rightarrow\pm\infty$, we will consider the boundary conditions that demand purely outgoing waves at both infinities. In other words, no waves are coming from either asymptotically flat region Konoplya:2005et ; Bronnikov:2021liv
@@
$R\sim e^{i\Omega r^{\ast}}\,\,\,\text{as}\,\,\,r^{\ast}\rightarrow\infty\,,$
@@
$R\sim e^{i\Omega r}r^{iM\omega^{2}/\Omega}\,\,\,\text{as}\,\,\,r\rightarrow\infty\,.$ (23)
Also, since the potential is symmetric about the throat of the wormhole, which is located at $r_{0}^{\ast}=0$, the solutions will be symmetric or anti-symmetric. Therefore, we impose the following boundary conditions at the throat, $\frac{dR}{dr^{\ast}}\left|{}_{r_{0}^{\ast}}=0\right.$ for the symmetric solutions and $R(r_{0}^{\ast})=0$ for the anti-symmetric solutions. These boundary conditions yields the even and odd overtones, respectively. It is worth mentioning that similar boundary conditions have been applied for other wormhole geometries, where the QNMs were obtained by direct integration of the wave equation Aneesh:2018hlp ; DuttaRoy:2019hij .
It is convenient to consider the following ansatz for the symmetric solutions of the radial equation (21) which satisfies the desired boundary conditions
$R(r)=e^{i\Omega r}r^{iM\omega^{2}/\Omega}\sum_{n=0}^{\infty}a_{n}\left(\frac{r-b_{0}}{r}\right)^{n}\,.$ (24)
@@
$c_{1}(n)$ $=$ $b_{0}^{3}\left(m^{2}-\omega^{2}\right)^{2}-ib_{0}^{2}\left(m^{2}-\omega^{2}\right)\left((4n+1)\sqrt{\omega^{2}-m^{2}}-2im^{2}M\right)+b_{0}\Bigg{(}m^{2}\bigg{(}l^{2}+l+n\left(8iM\sqrt{\omega^{2}-m^{2}}+5n+2\right)$
$+2iM\sqrt{\omega^{2}-m^{2}}+1\bigg{)}-\omega^{2}\left(l^{2}+l+n\left(4iM\sqrt{\omega^{2}-m^{2}}+5n+2\right)+iM\sqrt{\omega^{2}-m^{2}}+1\right)\Bigg{)}+$
$2M\left(-\left(l^{2}+l+n(7n+2)+1\right)\left(m^{2}-\omega^{2}\right)-4iMn\omega^{2}\sqrt{\omega^{2}-m^{2}}-iM\omega^{2}\sqrt{\omega^{2}-m^{2}}\right)\,,$
@@
$-6iM\sqrt{\omega^{2}-m^{2}}+8M^{2}\omega^{2}+1\Bigg{)}\bigg{)}+b_{0}^{2}\left(m^{2}-\omega^{2}\right)\left(2m^{2}M+i(2n-1)\sqrt{\omega^{2}-m^{2}}\right)+M\big{(}(2l(l+1)+n(18n-$
@@
$R(r)=(r-b_{0})^{1/2}e^{i\Omega r}r^{iM\omega^{2}/\Omega-1/2}\sum_{n=0}^{\infty}a_{n}\left(\frac{r-b_{0}}{r}\right)^{n}\,,$ (27)
@@
$c_{1}(n)$ $=$ $b_{0}\Bigg{(}m^{2}\left(4l(l+1)+4n\left(8iM\sqrt{\omega^{2}-m^{2}}+5n+7\right)+24iM\sqrt{\omega^{2}-m^{2}}+13\right)-\omega^{2}\bigg{(}4l(l+1)+$
@@
$4b_{0}^{3}\left(m^{2}-\omega^{2}\right)^{2}+2M\left(-(4l(l+1)+4n(7n+9)+15)\left(m^{2}-\omega^{2}\right)-16iMn\omega^{2}\sqrt{\omega^{2}-m^{2}}-12iM\omega^{2}\sqrt{\omega^{2}-m^{2}}\right)\,,$
@@
$2iM\sqrt{\omega^{2}-m^{2}}+12M^{2}\omega^{2}+1\bigg{)}+8b_{0}^{2}\left(m^{2}-\omega^{2}\right)\left(m^{2}M+in\sqrt{\omega^{2}-m^{2}}\right)+4M\bigg{(}\left(2l(l+1)+18n^{2}+n+3\right)$
@@
Then, after applying Gaussian elimination twice, the original five-term recurrence relations (IV.1) can be simplified to a three-term relation, taking on the following concise form
@@
The continued fraction must be truncated at some large index $N$ and the QNFs are obtained solving this equation numerically. In the following figures we show the behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for massive scalar field as a function of $mb_{0}$ using the CFM, for the fundamental mode, see Fig. 2, and 4, and the first overtone, see Fig. 3, and 5.
@@
Figure 2: The behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for massive scalar field as a function of $mb_{0}$ for the generalized Bronnikov-Ellis wormhole with $M/b_{0}=0.1$, $n=0$ (fundamental frequency), and $\ell=0,1,5$ using the CFM.
Figure 3: The behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for massive scalar field as a function of $mb_{0}$ for the generalized Bronnikov-Ellis wormhole with $M/b_{0}=0.1$, $n=1$ (first overtone), and $\ell=0,1,5$ using the CFM.
Figure 4: The behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for massive scalar field as a function of $mb_{0}$ for the generalized Bronnikov-Ellis wormhole with $M/b_{0}=0.25$, $n=0$ (fundamental frequency), and $\ell=0,1,5$ using the CFM.
Figure 5: The behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for massive scalar field as a function of $mb_{0}$ for the generalized Bronnikov-Ellis wormhole with $M/b_{0}=0.25$, $n=1$ (first overtone), and $\ell=0,1,5$ using the CFM.

### 10 Apr 2024 v3

# Quasinormal modes and bound states of massive scalar fields in wormhole spacetimes

@@ Abstract
In this work we explore the propagation of massive scalar fields on some wormhole backgrounds. On one side, we consider the Bronnikov-Ellis wormhole solution and wormhole geometries with a non-constant redshift function by introducing a gravitational mass $M$, which goes over into the Bronnikov-Ellis wormhole when the gravitational mass parameter vanishes. We employ the continued fraction method to calculate accurately the quasinormal frequencies of massive scalar fields, particularly focusing on low values of the angular number, and we show an anomalous behaviour of the decay rate of the quasinormal frequencies, for $n\geq\ell$. Also, we show that for a massive scalar field and $M\neq 0$ the effective potential allows potential wells for some values of the parameters which support bound states, which are obtained using the continued fraction method and they are characterized by having only a frequency of oscillation and they do not decay; however, for the Bronnikov-Ellis wormhole the effective potential do not support bound states. On the other side, we consider a wormhole geometry which is an exact solution of $f(R)$ modified gravity. For this geometry the quasinormal frequencies can be obtained analytically, being the longest-lived modes the ones with lowest angular number $\ell$. So, in this wormhole background the anomalous behaviour is avoided.
@@ I Introduction
Solutions of Einstein equations that connect different parts of the Universe or two different Universes can be found in General Relativity (GR) known as wormholes. The concept of the wormhole can be traced back to Flamm in 1916 Flamm and then wormhole-type solutions were found by Einstein and Rosen (ER), where they considered a physical space representing a physical space which was connected by a wormhole-type solution known as ER bridge Rosen . Then, the pioneering work of Misner and Wheeler MisWheel and Wheeler Wheel developed further the wormholes concept. Lorentzian wormholes were studied in GR Bronnikov:1973fh ; Ellis:1973yv ; Morris ; phantom ; Visser , in which conditions for traversable wormholes were found introducing a static spherically symmetric metric. However, a condition on the wormhole throat leads to the violation of the null energy condition (NEC). To allow the formation of traversable wormhole geometries a matter distribution of exotic or phantom matter has to be introduced phantom .
Recently there are many efforts to understand modifications of GR which are generated by the presence of a scalar field coupled to gravity. This coupling has important implications in local and global solutions in these theories known as scalar-tensor theories Fujii:2003pa . One of the best studied scalar-tensor theories is provided by the Horndeski Lagrangian Horndeski:1974wa . The Horndeski theory has been studied in short and large distances. In this theory black hole solutions were found Kolyvaris:2011fk ; Rinaldi:2012vy ; Kolyvaris:2013zfa ; Babichev:2013cya ; Charmousis:2014zaa , known as Galilean black holes, and also wormhole geometries were generated Korolev:2014hwa -Bakopoulos:2023tso .
@@
An important issue of the compact objects formed by black holes or wormholes is their stability. The relativistic collision of two of these compact objects produces gravitational waves which will help us to understand better the gravitational interaction. However, the recent LIGO detections Abbott:2016blz -TheLIGOScientific:2017qsa do not yet probe the detailed structure of spacetime beyond the photon sphere. Future observations may detect the ringdown phase, which is governed by a series of damped oscillatory modes at early times, which are known as quasinormal modes (QNMs) Vishveshwara:1970zz -Cardoso:2016rao . Future gravitational observations may provide information on different compact objects than black holes, objects without event horizons which are known as exotic compact objects (ECOs) Mazur:2001fv -Holdom:2016nek . Perturbations of black holes and wormholes were extensively studied in the literature. In the case that a test scalar perturbs black hole compact objects the partially reflected waves from the photon sphere (PS) mirror off the AdS boundary and re-perturb the PS to give rise to a damped beating pattern. In the case of a wormhole do not decay with time but have constant and equal amplitude to that of the initial ringdown. Therefore, the knowledge of QNMs and quasinormal frequencies (QNFs) can give us important information on the properties of compact objects and distinguish their nature. The QNMs give an infinite discrete spectrum which consists of complex frequencies, $\omega=\omega_{R}+i\omega_{I}$, where the real part $\omega_{R}$ determines the oscillation timescale of the modes, while the complex part $\omega_{I}$ determines their exponential decaying timescale (for a review on QNMs see Kokkotas:1999bd ; Berti:2009kk ).
@@
QNMs and QNFs have also been calculated in the background of wormholes. The perturbations of wormhole spacetime was investigated in Konoplya:2005et . It was found that the QNMs behave the same way as in black hole background and includes the quasinormal ringing and power-law asymptotically late time tails. However, in Konoplya:2016hmd it was found that the symmetric WHs do not have the same ringing behaviour of the BHs at a few various dominant multipoles. In the case of massive probe scalar field perturbations, it was shown Churilova:2019qph that wormholes with a constant redshift function do not allow for longest-lived modes, while wormholes with non-vanishing radial tidal force do allow for quasi-resonances. Calculating the high frequency QNMs in the case of the Morris-Thorne wormhole spacetime, the shape function of a spherically symmetric traversable Lorenzian wormhole near its throat was constructed in Konoplya:2018ala . Scalar and axial QNMs of massive static phantom WHs were calculated in Blazquez-Salcedo:2018ipc .
@@
As we already discussed in the case of the Horndeski scalar-tensor theory if matter, parameterized by a phantom scalar field, is strongly coupled to gravity then bound states are formed Koutsoumbas:2018gbd ; Chatzifotis:2020oqr . In this work we follow a much simpler approach. We consider the Bronnikov-Ellis wormhole solution Bronnikov:1973fh ; Ellis:1973yv and wormhole geometries with non-constant redshift function by introducing a gravitational mass $M$, which goes over into the Bronnikov-Ellis wormhole when $M=0$. We perturb the massive wormhole background by a test massive scalar field. Using the continued fraction method we calculate accurately the QNFs for low values of the angular number. We find stability of the background wormhole and calculating the effective potential we show that for an interplay of the mass parameter and the phantom scalar charge, potential wells are generated supporting bound states.
We also study the QNFs in a wormhole spacetime supported by a phantom scalar field in which the scalar curvature is generalized to an $f(R)$ function Karakasis:2021lnq ; Karakasis:2021rpn ; Tang:2019qiy ; DeFalco:2021ksd ; Karakasis:2021tqx . In this model the scalar field self-interacts with a mass term potential which is derived from the scalar equation and in the resulting $f(R)$ model the scalar curvature is modified by the presence of the scalar field and it is free of ghosts and avoids the tachyonic instability.
The work is organized as follows. In Section II we review the solutions of the traversable wormholes. In Section III we study the perturbations of massive scalar field in the wormhole geometry. In Section IV we calculate the QNFs of the non-constant redshift function wormholes, resulting in a non-zero tidal radial force, and the Bronnikov-Ellis wormhole numerically using the continued fraction method. Also, we show that bound states are allowed by the wormholes with non-vanishing radial tidal force and the frequencies of the lowest bound states are calculated using the continued fraction method. In Section V we calculate the QNFs of an $f(R)$ gravity wormhole analytically and using the WKB method. In Section VI are our conclusions.
@@ II Review of traversable wormholes
To ensure the absence of horizons and singularities requires $\Phi(r)\neq 0$ which means that $\Phi(r)$ is finite throughout the spacetime. The Ellis Ellis:1973yv is a wormhole solution of an action that consists of a pure Einstein-Hilbert term and a scalar field with negative kinetic energy
@@
Given the metric ansatz (5) and setting $V(r)=0$, the aforementioned equations yield the following solution
$\Phi(r)$ $=$ $0\,,$ (8)
@@
which corresponds to the Bronnikov-Ellis wormhole in the limit $M\rightarrow 0$ Bronnikov:1973fh ; Ellis:1973yv . Note that the parameter $b_{0}$ specifies the charge of the phantom scalar field.
@@
For the wormhole geometry (13), the obtained expressions are as follows
@@
III Massive scalar field perturbations
@@
$\Psi(t,r,\theta,\varphi)=e^{-i\omega t}\>\frac{u(r)}{r}\>Y_{\ell}(\Omega)\,.$ (16)
@@
$\frac{\mathrm{d}^{2}u}{\mathrm{d}r^{\ast}{}^{2}}+[\omega^{2}-V(r^{\ast})]u=0\,,$ (17)
@@
IV Bronnikov-Ellis wormholes and non-constant redshift function wormholes
@@
which is plotted in Fig. 1, where a barrier-like shape is observed. Additionally, a well may form near the throat for a massive scalar field, $M\neq 0$ and at low values of $\ell$, while for high values of $\ell$, the well disappears; therefore, the potential can support bound states for low values of $\ell$.
Figure 1: Effective potential of the wormhole spacetime (13), for $M/b_{0}=0.25$, $mb_{0}=0.8$ and $\ell=0,1,2$.
For $M=0$ the effective potential transforms into $V(r)=q^{2}/r^{4}+\ell(\ell+1)/r^{2}+m^{2}$, and it is straightforward to demonstrate that it cannot give rise to a well; hence, bound states are prohibited in this case.
@@
$(2M-r)\left(b_{0}^{2}-r^{2}\right)u^{\prime\prime}(r)+\frac{\left(-3Mb_{0}^{2}+Mr^{2}+b_{0}^{2}r\right)}{r}u^{\prime}(r)+$
$\frac{M\left(\left(2m^{2}r^{2}+2\ell^{2}+2\ell-1\right)r^{2}+3b_{0}^{2}\right)-r\left(r^{2}\left(\ell^{2}+\ell+r^{2}(m^{2}-\omega^{2})\right)+b_{0}^{2}\right)}{r^{2}}u(r)=0\,,$ (22)
which we will solve numerically to determine both the QNFs and the frequencies of bound states.
@@ IV.1 Quasinormal frequencies
In this section we use the continued fraction method (CFM) to calculate accurately the QNFs for massive scalar fields propagating on the Bronnikov-Ellis wormhole and non-constant redshift function wormholes described by Eq. (13). The CFM was devised by Leaver to compute the QNMs of Schwarzschild and Kerr black holes Leaver:1985ax ; Leaver:1990zz , and improved later by Nollert Nollert:1993zz . It has been used to computed the QNMs in several black holes backgrounds, see for instance Cardoso:2003qd ; Konoplya:2013rxa ; Richartz:2014jla ; Chowdhury:2018pre ; Becar:2022wcj . However, to our knowledge this is the first time that this method is applied to calculate the QNFs in wormhole geometries.
To solve the radial equation it is necessary to specify the boundary conditions. Given the asymptotically flat nature of the wormhole geometries and the fact that the effective potential forms a barrier approaching $V(r^{\ast})\rightarrow m^{2}$ as $r^{\ast}\rightarrow\pm\infty$, we will consider the boundary conditions that demand purely outgoing waves at both infinities. In other words, no waves are coming from either asymptotically flat region Konoplya:2005et ; Bronnikov:2021liv
@@
$u\sim e^{i\Omega r^{\ast}}\,\,\,\text{as}\,\,\,r^{\ast}\rightarrow\infty\,,$
@@
$u\sim e^{i\Omega r}r^{iM\omega^{2}/\Omega}\,\,\,\text{as}\,\,\,r\rightarrow\infty\,.$ (24)
Also, since the potential is symmetric about the throat of the wormhole, which is located at $r_{0}^{\ast}=0$, the solutions will be symmetric or anti-symmetric. Therefore, we impose the following boundary conditions at the throat, $\frac{du}{dr^{\ast}}\left|{}_{r_{0}^{\ast}}=0\right.$ for the symmetric solutions and $u(r_{0}^{\ast})=0$ for the anti-symmetric solutions. These boundary conditions yields the even and odd overtones, respectively. It is worth mentioning that similar boundary conditions have been applied for other wormhole geometries, where the QNMs were obtained by direct integration of the wave equation Aneesh:2018hlp ; DuttaRoy:2019hij .
It is convenient to consider the following ansatz for the symmetric solutions of the radial equation (IV) which satisfies the desired boundary conditions
$u(r)=e^{i\Omega r}r^{iM\omega^{2}/\Omega}\sum_{n=0}^{\infty}a_{n}\left(\frac{r-b_{0}}{r}\right)^{n}\,.$ (25)
@@
$c_{1}(n)$ $=$ $b_{0}^{3}\left(m^{2}-\omega^{2}\right)^{2}-ib_{0}^{2}\left(m^{2}-\omega^{2}\right)\left((4n+1)\sqrt{\omega^{2}-m^{2}}-2im^{2}M\right)+b_{0}\Bigg{(}m^{2}\bigg{(}\ell^{2}+\ell+n\left(8iM\sqrt{\omega^{2}-m^{2}}+5n+2\right)$
$+2iM\sqrt{\omega^{2}-m^{2}}+1\bigg{)}-\omega^{2}\left(\ell^{2}+\ell+n\left(4iM\sqrt{\omega^{2}-m^{2}}+5n+2\right)+iM\sqrt{\omega^{2}-m^{2}}+1\right)\Bigg{)}+$
$2M\left(-\left(\ell^{2}+\ell+n(7n+2)+1\right)\left(m^{2}-\omega^{2}\right)-4iMn\omega^{2}\sqrt{\omega^{2}-m^{2}}-iM\omega^{2}\sqrt{\omega^{2}-m^{2}}\right)\,,$
@@
$-6iM\sqrt{\omega^{2}-m^{2}}+8M^{2}\omega^{2}+1\Bigg{)}\bigg{)}+b_{0}^{2}\left(m^{2}-\omega^{2}\right)\left(2m^{2}M+i(2n-1)\sqrt{\omega^{2}-m^{2}}\right)+M\big{(}(2\ell(\ell+1)+n(18n-$
@@
$u(r)=(r-b_{0})^{1/2}e^{i\Omega r}r^{iM\omega^{2}/\Omega-1/2}\sum_{n=0}^{\infty}a_{n}\left(\frac{r-b_{0}}{r}\right)^{n}\,,$ (28)
@@
$c_{1}(n)$ $=$ $b_{0}\Bigg{(}m^{2}\left(4\ell(\ell+1)+4n\left(8iM\sqrt{\omega^{2}-m^{2}}+5n+7\right)+24iM\sqrt{\omega^{2}-m^{2}}+13\right)-\omega^{2}\bigg{(}4\ell(\ell+1)+$
@@
$4b_{0}^{3}\left(m^{2}-\omega^{2}\right)^{2}+2M\left(-(4\ell(\ell+1)+4n(7n+9)+15)\left(m^{2}-\omega^{2}\right)-16iMn\omega^{2}\sqrt{\omega^{2}-m^{2}}-12iM\omega^{2}\sqrt{\omega^{2}-m^{2}}\right)\,,$
@@
$2iM\sqrt{\omega^{2}-m^{2}}+12M^{2}\omega^{2}+1\bigg{)}+8b_{0}^{2}\left(m^{2}-\omega^{2}\right)\left(m^{2}M+in\sqrt{\omega^{2}-m^{2}}\right)+4M\bigg{(}\left(2\ell(\ell+1)+18n^{2}+n+3\right)$
@@
For $M=0$, the coefficients $c_{4}(n)$ become zero for both symmetric and antisymmetric solutions; hence, the expansion coefficients satisfy four-term recurrence relations for the Bronnikov-Ellis wormhole. Then, after applying Gaussian elimination twice for $M\neq 0$, and once for $M=0$, the original recurrence relations (IV.1) can be simplified to a three-term relation, taking on the following concise form
@@
The continued fraction must be truncated at some large index $N$ and the QNFs are obtained solving this equation numerically. In the following figures we show the behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for massive scalar field as a function of $mb_{0}$ using the CFM, for the fundamental mode, see Figs. 2, 4 and 6, and the first overtone, see Figs. 3, 5 and 7.
@@
Figure 2: The behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for massive scalar field as a function of $mb_{0}$ for the Bronnikov-Ellis wormhole with $n=0$ (fundamental frequency), and $\ell=0,1,5$ using the CFM.
Figure 3: The behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for massive scalar field as a function of $mb_{0}$ for the Bronnikov-Ellis wormhole with $n=1$ (first overtone), and $\ell=0,1,5$ using the CFM.
Figure 4: The behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for a massive scalar field as a function of $mb_{0}$ for the wormhole spacetime described by Eq. (13), with $M/b_{0}=0.1$, $n=0$ (fundamental frequency), and $\ell=0,1,5$ using the CFM.
Figure 5: The behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for a massive scalar field as a function of $mb_{0}$ for the wormhole spacetime described by Eq. (13), with $M/b_{0}=0.1$, $n=1$ (first overtone), and $\ell=0,1,5$ using the CFM.
Figure 6: The behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for a massive scalar field as a function of $mb_{0}$ for the wormhole spacetime described by Eq. (13), with $M/b_{0}=0.25$, $n=0$ (fundamental frequency), and $\ell=0,1,5$ using the CFM.
Figure 7: The behaviour of $Re(\omega)b_{0}$ and $-Im(\omega)b_{0}$ for a massive scalar field as a function of $mb_{0}$ for the wormhole spacetime described by Eq. (13), with $M/b_{0}=0.25$, $n=1$ (first overtone), and $\ell=0,1,5$ using the CFM.